Crew 219 Operations Report 11-JAN-2020
Name of person filing report: Alejandro Perez
Non-nominal systems: N/A
Hours run: 14.5
From what time last night: 17:00
To what time this morning: 07:30
List of any additional daytime hours when the generator was run: N/A
Solar- SOC% (Before the generator is run at night): 77%
08:40 = 96%; 09:30 = 94%; 10:20 = 92%; 11:00 = 90%; 11:30= 89%; 12:00=
89%; 12:30= 88%; 13:00 = 85%; 13:30= 85%; 15:00= 82%; 16:00= 80%;
Diesel Reading – 65%
Station Propane Reading- 45%
Ethanol Free Gasoline- N/A
Water (loft tank) (gal) – 50
Water Meter (units)- 0147567,5
Water (static tank) (gal) – 444
Static to loft Pump used – Yes
Water in Green Hab (gal): 179.45
Water in Science Dome (gal): 0
Toilet Tank Emptied – Yes
Diemos rover used: Still in the Work shop
Sojourner rover used: Assigned to Director
Spirit Rover used- Yes
Starting Hours: 124.5, 124.5
Beginning Charge: 100%, 100%
Ending Hours: 124.5, 124.9
Ending Charge: 23%, 47%
Opportunity Rover used: Still in workshop
Curiosity Rover used: Yes
Starting Hours: 129.9
Beginning Charge: 100%
Ending Hours: 130.3
Ending Charge: 39%
Notes on Rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use:
Oil Added? No
# hours the ATV’s were used today:
Notes on ATVs:
Habcar used and why, where?
CrewCar used and why, where?
General Notes and Comments:
Summary of the internet: Nominal
Summary of Suit and Radios: Suit #5 was tested today and worked properly.
Summary of Hab: Air filter for Heater has been replaced with a Clean one. New Teflon tape was used on the Loft tank feedline.
Summary of Science Dome operations: Temperatures are hovering around 7°C during the day.
Summary of Ram operations: Nominal
Summary of any observatory issues: Nominal
Summary of Health and safety issues: Nominal
Questions, concerns, and requests to Mission Support: Metal dish scrubber and a brush to wash dishes for pots and pans.